OpenStudy (anonymous):

Determine whether the table below represents an arithmetic sequence. If so, write the formula. X-1,2,3,4,5 Y-1,-8,-17,-26,-35

OpenStudy (anonymous):

First One: X is an Arithmetic Progression with common difference of 1 a = 1 and d = 1 tn = a + (n - 1)*d = 1 n - 1 = n tn = n Second One: d = -8 - 1 = -9 d = -17 + 8 = -9 Y is also an Arithmetic Progression with common difference of (-9) a = 1 and d = -9 tn = 1 + (n - 1)*(-9) tn = 1 -9n + 9 tn = 10 - 9n